What are the magnitude and direction of the electric field that will balance the


weight of (a) an electron and (b) a proton?


1
Expert's answer
2021-12-02T10:04:53-0500

"qE=mg,\implies E=\frac{mg}q,"

a)

"E_1=\frac{m_1g}{q_1}=-5.58\cdot 10^{-11}~\frac NC,"

b)

"E_2=\frac{m_2g}{q_2}=1.02\cdot 10^{-7}~\frac NC."